if f (x)=x^2-18x+13 find and simiplify f(x+h)-f(x)/h h=0

Answer by stanbon(75887)

if f(x)=x^2-18x+13 find and simiplify f(x+h)-f(x)/h h=0
-------------
f(x+h) = (x+h)^2-18(x+h)+13
f(x+h) = x^2+2hx+h^2-18x-18h+13
-----
Subtract x^2-18x+13 to get:
2hx+h^2-18h
----
Divide by "h": to get:
2x +h - 18
================
Cheers,
Stan H.
=====================